A 145 mH inductor with 71 Ω resistance is connected in series to a 21 μF capacitor and a 60 Hz , Vrms = 60 Vsource.

Part A= Calculate the rms current.
Answer #1

XL = 2 pi f L = 2 pi * 60 * 145 * 10-3 = 54.66 ohm

XC = 1 / 2 pi f C = 1 / (2 pi * 60 * 21 * 10-6) = 126.3 ohm

Z = sqrt [R2 + (XL - XC)2] = sqrt [712 + (54.66 - 126.3)2] = 100.8 ohm

Vrms = I Z

I = 60 / 100.8

rms current = 0.594 A